# Polycubes via Dual Loops

### v1

@@ 5 Polycube segmentation construction
Our characterization of polycubes gives rise to an iterative algorithm to construct a polycube segmentation of a given input surface $\mathcal{S}$ and, by extension, a polycube map and a hexahedral mesh. In the following we first briefly sketch the steps that our proof-of-concept algorithm and its implementation take and then showcase our results. Note that neither algorithm nor implementation have been optimized for speed or quality; they simply serve to establish the feasibility of our approach.
We initialize our iterative algorithm with a single cube for genus-0 surfaces or a suitable polycube of genus $g$ for surfaces of genus $g$. In the future we plan to develop an algorithm that automatically constructs suitable initial polycubes and corresponding loop structures for an input surface $\mathcal{S}$; for our proof-of-concept we used manually-crafted polycubes and loop structures. To embed the initial loop structure on the input surface $\mathcal{S}$ we follow Campen and Kobbelt [5] and use shortest paths in a suitably weighted graph on $\mathcal{S}$. Here we can restrict ourselves to those parts of the graph which intersect specific loop segments (see Section 4.1). We attempt to embed loops to align with their label and orientation.
We now iteratively add or remove loops. To evaluate the quality of a loop structure we need to primalize it into a polycube segmentation. We first place one polycube corner into each loop region, often on or close to a critical point, since that tends to reduce distortion. We also attempt to align polycube corners within the same zone, since this results in segmentation patches which are more rectangular. Then we find non-intersecting paths on $\mathcal{S}$ to connect adjacent polycube corners, guided by the dual loops embedded on $\mathcal{S}$.
There are often many different loops that can be added to or removed from a valid loop structure. We use a simple evolutionary algorithm to choose loops that improve the polycube segmentation. This algorithm uses mutation and selection only (no crossover); every candidate solution is a valid loop structure. To asses the quality of a polycube segmentation we use a simply metric that approximates the distortion of the corresponding polycube map: the alignment of each surface patch with its assigned axis. Figure 12 shows some of our results for surfaces genus 0 and 1.
@@ 6 Future Work
The results of our proof-of-concept algorithm are very promising and show that our polycube characterization via dual loops, as well as the modification operations, can be efficient and effective in practice. However, various avenues for future work remain.
On the theoretical side we plan to establish which operations are necessary to construct any loop structure. We are currently not aware of a polycube of genus 0 that cannot be constructed via loop addition and removal. For polycubes of genus $g>1$ we also do not seem to require other operations while moving within the set of polycubes of genus $g$. However, loop addition and removal cannot re-orient holes or change the genus.
On the practical side, we would like to develop an algorithm that constructs an initial polycube that captures global features of input surfaces of any genus. Furthermore, an optimal embedding of the loop structure on the surface should ideally align with salient features, such a ridges, valleys or handles. Finally, we would like to develop a method that finds an optimal primalization of our embedded loop structures, where optimality will necessarily depend on the use case.

### v3

@@ 5 Polycube segmentation construction
Our characterization of polycubes gives rise to an iterative algorithm to construct a polycube segmentation of a given input surface $\mathcal{S}$ and, by extension, a polycube map and a hexahedral mesh. In the following, we first briefly sketch the steps that our proof-of-concept algorithm and its implementation take and then showcase our results. Note that neither algorithm nor implementation have been optimized for speed or quality; they simply serve to establish the feasibility of our approach.
We initialize our iterative algorithm with a single cube for genus-0 surfaces or a suitable polycube of genus $g$ for surfaces of genus $g$. In the future, we plan to develop an algorithm that automatically constructs suitable initial polycubes and corresponding loop structures for an input surface $\mathcal{S}$; for our proof-of-concept, we used manually-crafted polycubes and loop structures. To embed the initial loop structure on the input surface $\mathcal{S}$, we follow Campen and Kobbelt [5] and use shortest paths in a suitably weighted graph on $\mathcal{S}$. Here, we can restrict ourselves to those parts of the graph that intersect specific loop segments (see Section 4.1). We attempt to embed loops to align with their label and orientation.
We now iteratively add or remove loops. Figure 12 illustrates the steps of our algorithm through a series of snapshots. It begins with an initial polycube dual loop structure for a genus-1 model, followed by successive steps adding X-loops, Y-loops, and Z-loops. To evaluate the quality of a loop structure we need to primalize it into a polycube segmentation. We first place a polycube corner into each loop region while attempting to align corners within the same zone, to create segmentation patches that are mostly rectangular. Then, we find non-intersecting paths on $\mathcal{S}$ to connect adjacent polycube corners, guided by the dual loops embedded on $\mathcal{S}$.
There are often many different loops that can be added to or removed from a valid loop structure. We use a simple evolutionary algorithm to choose loops that improve the polycube segmentation. This algorithm uses mutation and selection only (no crossover); every candidate solution is a valid loop structure. To asses the quality of a solution, we use a metric that approximates the distortion of the corresponding polycube map: the alignment of each surface patch with its assigned axis. Figure 13 shows some of our results for various models with varying complexity and genus.
@@ 6 Future Work
Our proof-of-concept algorithm demonstrates the potential of our method. While the current implementation is minimal and lacks extensive experimental evaluation, it nevertheless provides a glimpse of the future possibilities. Various avenues for future work remain.
On the theoretical side, we plan to establish which operations are necessary to construct any loop structure. We are currently not aware of a polycube of genus 0 that cannot be constructed via loop addition and removal. For polycubes of genus $g>1$, we also do not seem to require other operations while moving within the set of polycubes of genus $g$. However, loop addition and removal cannot re-orient holes or change the genus. Our current method constructs polycube maps which only map the surface of the polycube to the input surface. It would be interesting to see if and how our method extends to volumetric maps.
Figure 13: Embedded polycube loop structure, the polycube segmentation, and the corresponding polycube constructed by our proof-of-concept implementation.
On the practical side, we would like to develop an algorithm that constructs an initial polycube that captures global features of input surfaces of any genus. Furthermore, an optimal embedding of the loop structure on the surface should ideally align with salient features, such as ridges, valleys or handles. Finally, we would like to develop a method that finds an optimal primalization of our embedded loop structures, where optimality will necessarily depend on the use case.